Fugacity in terms of compressibility factor and molar volume

Starting from the fundamental expression of fugacity [dG = RT d(ln f)], derive a general expression for the fugacity coefficient (φ) in terms of the compressibility factor (Z) and molar volume (V).
Answer #1

* Solution :- PAGE 1 Considering the conditions of palometers from the question 2- 8.3/4 Imolly TEOC = 273K H-10 C-12W N = 14NACE-2 0 O = VA S V; RT (b) du when = Rr la Cue-b J - 9so themel compression Wyob 3 i) 2+0) js łatbp P ale Iola OT 2 P (2T+b)PAGE-3 E (т-с), к?кт {ccla+20) xed в (- ۴ می ф, er(i=a); RT ра (o-b) put in eyes, In d2 = { [ er Po-l] SP [Intra frip •ps:E 1

answered by: ANURANJAN SARSAM